Scores in a standardized exam are normally distributed with mean μ = 510 and standard deviation σ = 60. The middle 75% of scores are classified as “average”. Find the range of scores that are average according to this definition.


1
Expert's answer
2021-05-07T11:43:57-0400

"\mu=510,\sigma=60"


Middle 75% are classified as avergae-

"\alpha=0.25"


"z_{\frac{\alpha}{2}}=z_{0.125}=1.1503"


Confidence Interval "= \mu\pm z_{0.125}\times 60"

"=510\pm 1.1503\times 60\\=(510\pm 69.018)\\=(440.98,579.018)"